# Improving Student-AI Interaction Through Pedagogical Prompting: An Example in Computer Science Education

### v1

@@ 5.3. A Step-by-Step Pedagogical Prompt Builder
To effectively scaffold students in creating pedagogical prompts, we divided the prompt construction process into six steps based on the six main components of a pedagogical prompt described in Section 3, according to the segmenting principle to reduce unnecessary mental effort (Mayer and Pilegard, 2005). For the Problem Context step, the builder lists key information that students must explicitly incorporate into their final prompt and presents it as a short-answer question. Other steps begin with a targeted multiple choice question (Figure 8-C1). After selecting or reading the corresponding step description, participants construct and submit a draft segment of the prompt (Figure 8-C2). When writing a prompt segment, students see their previously completed segments displayed before the current text input area as contextual reminders (Figure 7). Students then submit their prompt segment to receive immediate, AI-generated feedback (Figure 8-C3). We defined specific acceptance criteria for each step of the prompt-building process, and employed OpenAI’s GPT-4o (specifically gpt-4o-2024-08-06) to provide personalized, elaborated actionable feedback, highlighting any omissions or areas for improvement. If the student cannot pass all the criteria after three attempts, a sample solution for this step will be provided (Figure 8-C4) to help the student finish the current prompting step (Figure 8-C5). Upon finishing constructing a prompt (Figure 7), students are encouraged to use this prompt to interact with AI, such as putting it into their Chatbot of choice (e.g., ChatGPT, Claude) and observing the AI response.
@@
Figure 6. Previous prompt segments. Students will be able to see prompt segments that they write as previous steps above the input field above the current input field.
Figure 7. A full prompt review. Students will be able to see the entire prompt themselves built after the step-by-step guidance, after completing the current problem scenario.
6. User Study: Understand Student Learning and Perceptions
@@
Figure 8. Example of completing Problem Identification step in the prompt builder workflow. The student first (C1) considers and selects the most accurate option within the given list, and then reaches the (C2) prompt writing task, in which they are required to complete the current snippet of the prompt given their previous selection. When the student has drafted a prompt, (C3) an LLM-based feedback system will provide personalized feedback and provide the student with actionable suggestions. If the student still has difficulties passing the prompt validator, (C4) a sample solution will be shown to the student, and they have a chance to review and paste it into the answer box. (C5) When a students= have successfully completed the prompt writing task for the current step, they will be directed to the next step.
@@
To understand the effectiveness of this learning intervention, we conducted a evaluation study with 22 students with IRB (Institutional Review Boards) approval. The RQs were:
@@
6.1. Participants & Study Procedure
@@
At the beginning of each session, participants first completed a consent with a pre-screening survey that included questions about their programming learning level, self-reported familiarity with AI and generative AI tools. Then they got to a pre-test that included 9 questions. The learning session started with a detailed on-boarding demo video. After that, the participant completed three scenario-based learning tasks using our system in randomized order to decrease the ordering effect. Then they completed a post-test with isomorphic questions as the pre-test. After the post-test, they completed a post-survey containing likert questions and open-ended reflection questions to reflect on their learning and interaction with the system. All the sections were not time-limited, participants were allowed to learn and finish the tests at their own pace. Each participant received a \$30 per hour CAD Gift Card after the study. The detailed procedure can be found in Figure 9.
More specifically, the pre-test and post-test each included nine questions in two types. Students first finished three scenario-based prompt writing tasks with a student-selected commercial AI chatbot, with ChatGPT as the default option. The scenarios were isomorphic across the pre-test, post-test, and learning session. Then students completed six open-ended reflection questions (two for each scenario) based on their responses to the corresponding prompt-writing tasks. Below is an example isomorphic scenario (Scenario 1), along with its corresponding prompt-writing task and reflection questions (Figure 10).
@@
Figure 9. User study procedure. Pre-test: participants spent 25 minutes to complete a survey, 3 prompting questions, and 3 reflection questions on the previous 3 prompts; learning session: participants began with watching a 2-3 minutes demo video of the system, and them spent 30-40 minutes to write pedagogical prompts for 3 different scenarios with scaffolding; post-test: participants completed 3 prompting questions, 3 reflection questions, and a post-survey in order.
@@
Figure 10. An example isomorphic scenario in pre-test with its corresponding prompt-writing task and reflection questions
6.2. Data Analysis
@@
To answer RQ1, two graders manually graded student written prompts in the pre-test and post-test using the rubric extracted from the detailed definition (Section 5.3.1) of a high-quality pedagogical prompt. Specifically, two graders independently graded 15% of the data, half randomly selected from the pre-test and half from the post-test. The graders reached an 85% agreement in all fields (Cohen Kappa=0.72 - high agreement in inter-rater reliability (McHugh, 2012)) and further refine the grading rubric. With the refined rubric, one grader graded the remaining data. For RQ2, RQ3 and RQ4, we first summarized and reported students’ quantitative survey results, then uncovered the details based on their open-ended answers.
7. User Study Results
@@
7.1. RQ1: Students’ ability to write pedagogical prompts significantly improved after the learning intervention.
@@
For each component of a pedagogical prompt, we calculated students’ learning gain using post-test - pre-test. Then we applied Wilcoxon signed-rank tests (Gehan, 1965) because student test data is not normally distributed. Results show students’ (N=22) significant learning gains from pre-test to post-test for ALL 6 intended components (p ¡ .001 (see Table 1), indicating the initial success of this learning intervention in teaching pedagogical prompt.
@@
Figure 11. P5’s written prompt to the pre-test question and its corresponding isomorphic post-test question.
Table 1. Student test performance and learning gain for six pedagogical prompt components
@@
Tutoring Protocols 0.44 0.0 1.0 0.83 1.0 0.0 0.39 0.5 1.0 100.00 ¡.001*** 0.82
Difficulty Identification 0.70 1.0 1.0 0.95 1.0 0.0 0.26 0.0 0.8 0.00$\dagger$ ¡.001*** 0.86
AI’s Role 0.00 0.0 0.0 0.62 1.0 1.0 0.62 1.0 1.0 0.00$\dagger$ ¡.001*** 0.86
Problem Context 0.77 1.0 0.5 0.94 1.0 0.0 0.17 0.0 0.5 62.00 ¡.001*** 0.84
Learner’s Level 0.08 0.0 0.0 0.65 1.0 0.9 0.58 1.0 1.0 36.50 ¡.001*** 0.85
Guardrails 0.25 0.0 0.8 0.82 1.0 0.0 0.56 1.0 1.0 42.00 ¡.001*** 0.85
  • * $p$05, ** $p$01, *** $p$001. Wilcoxon signed-rank test used for non-parametric data; interquartile range (IQR) is reported to describe variability in scores.
  • $\dagger$ For Difficulty Identification and AI’s Role, a Wilcoxon test yielded $z=0.00$, indicating that all non-zero changes were positive after excluding ties.
7.2. RQ2: Students perceptions of their ability to effectively use LLM for learning have been corrected and increased after the learning intervention.
Students reported relatively high confidence in their ability to use LLMs effectively for learning in the pre-test (Figure 12, first row; M = 3.72 of 5). However, their self-reported confidence did not accurately reflect their actual prompting ability. Specifically, if students had an accurate self-assessment of their skills, a significant positive correlation between confidence and pre-test performance would be expected (Baxter and Norman, 2011). In contrast, in our result, a Pearson correlation showed a negative but not significant relationship between the perceived confidence score and pre-test performance (r = –.15, p = .498), suggesting a misalignment between the perception and actual abilities of the learners before the intervention. This pattern aligns with findings from Prather et al. (2024), suggesting that students may hold a false sense of confidence in their LLM-based learning, which can potentially lead to conceptual gaps and falling behind in course content. After the learning session, participants’ perceived confidence increased to an average of 4.63 of 5 (Figure 12-row 2), reflecting both greater perceived confidence from pre-survey to post-survey (p = .001, z = 13.00, effect size = .81, Wilcoxon signed-rank test) and a significant positive alignment between their perceived ability and actual performance (r = .54, p = .032, Pearson correlation). Furthermore, as indicated in Figure 12 (row 3 to row 4), students reported perceived improvements in identifying their specific struggles, although the improvement is not significant (p = .071, z = 18.00, Wilcoxon signed-rank test).
@@
Figure 12. Changes in students’ self-reported ability to use LLM tools for learning before and after the learning intervention
Figure 13. Students’ willingness to use pedagogical prompting in future learning and perceptions of the system experience reported in the post-survey
7.3. RQ3: Most students are extremely willing to apply pedagogical prompts in their future learning.
@@
7.4. RQ4: Students rated the system highly, despite occasional moments of frustration.
@@ 8. Discussion
8.1. Instructors: need for an interactive system that delivers brief instructional activities to teach pedagogical prompting for all CS novices
@@
8.1.1. Instructors prefer a brief but focused pedagogical prompting instruction session
@@
8.1.2. Pedagogical prompting should be taught to all students in CS classes, regardless of their major and level of programming.
There is a consensus on the targeted student population. Instructors think both CS majors and non-majors should learn pedagogical prompting, regardless of their major or level of programming. Recent studies support providing AI education to this broad audience (Shernoff, 2024; Walter, 2024; Denny et al., 2024). For instance, Walter emphasizes the importance of AI literacy and prompt engineering for all students regardless of their major programs, highlighting their role in fostering critical thinking skills essential in the age of AI (Walter, 2024). Similarly, Denny et al. discuss the transformative impact of generative AI on computing education, underscoring the need for students across disciplines to engage with these tools thoughtfully (Denny et al., 2024). These findings suggest that equipping all students with the skills to use AI tools pedagogically is crucial. It’s not only for students’ own academic success, but also for their preparedness in a workforce increasingly influenced by AI technologies.
@@
8.1.3. An interactive instructional system is the most preferred way for delivering the training
Regarding the medium of delivering pedagogical prompting instruction, an interactive system is the top choice. This finding demonstrates instructors’ preference for a more engaging instructional method. Their preferred ways align well with the ICAP framework (Chi and Wylie, 2014), which categorizes learning activities by the level of cognitive engagement they promote. Interactive instructional systems encourage students to either construct prompts themselves or actively select and refine them through system-guided scaffolding, both of which fall into the Constructive or Interactive modes of engagement. These levels are shown to produce deeper learning outcomes because learners generate new knowledge or meaning beyond the information presented. In contrast, alternatives such as demo videos or static spreadsheets primarily foster Passive or at best Active engagement, where students receive or manipulate information without generating new understanding. Practice questions can also be a viable option; however, compared to an interactive system, they may lack personalized, immediate feedback (Shute, 2008), as well as autonomy, which could induce boredom according to Pekrun and Perry’s Control-Value Theory of Achievement Emotions (Pekrun and Perry, 2014).
8.1.4. Selecting vs. Writing: Designing Interactions to Balance Efficiency and Active Skill Development
@@ 8.1.3. An interactive instructional system is the most preferred way for delivering the training
8.2. Learners: Developing Cognitive and Metacognitive Skills to Strengthen Learner Agency in the Age of Generative AI
@@
8.2.1. Prompting skill is a pressing need for nowadays learners, and scaffolded writing practice is an effective, immediate next-step that can significantly increase learners’ prompting ability in using AI for learning
@@
As generative AI becomes an inevitable component of future learning environments, it is crucial to prepare students to engage with it effectively rather than passively or improperly (Prather et al., 2025). Yet, learners, especially novices, often lack the strategies for effective AI-based help-seeking (Xiao et al., 2024b), and the misuses of AI can negatively impact learning outcomes (Bastani et al., 2024). Therefore, equipping students with the skills to use AI tools purposefully and critically will be an essential immediate next-step in education.
@@
Our system provides a solution to achieve this instructional goal in prompting learning in a short time period, as indicated by the promising results of the user study. After learning with our system, both learners’ ability and perceived confidence in pedagogical prompting significantly improved from pre-test to post-test. As shown in Table 1, participants made the most substantial gains in articulating the Learning Method and specifying the AI’s Role, which are critical components of effective pedagogical prompting. These improvements suggest that learners became more capable of providing structured, instructional guidance to the AI, thus enhancing the overall quality and clarity of their prompts. Mapping these gains to Nelson-Le Gall’s five-stage help-seeking model, we can observe strong developments in Stage 3: Identifying a potential helper, and Stage 4: Employing strategies to elicit help. By learning how to clearly specify the learning method and contextualize the AI’s role, students demonstrated an improved ability to recognize the AI as a valid and appropriate source of help for learning tasks. More importantly, they learned to strategically phrase their prompts in ways that actively engage the AI in a pedagogically useful manner.
@@
This success may be attributed to the system’s ability to stimulate learners’ psychological engagement (Stull and Mayer, 2007) through repeated practice (Pedersen et al., 2016) on constructive-level tasks (Chi and Wylie, 2014). We deliberately designed an interaction-heavy system, recognizing that while many existing works assume a high degree of learner self-regulation, this assumption often does not hold in real-world educational settings, as learners frequently demonstrate unregulated or suboptimal learning behaviors (Jansen et al., 2020; Molenaar et al., 2023; Xiao et al., 2024b). To address this, we embedded interactions that involve productive struggles (Gresalfi et al., 2009) and are difficult to ”game,” encouraging meaningful engagement without prerequisites on learners’ self-regulation. Since the effectiveness of the learning experience has been demonstrated, a next-step for future research is to investigate the system’s learning efficiency by optimizing learning gains while minimizing potential redundancy and learner frustration.
@@
8.2.2. Meta-cognitive skills can further strengthen students’ ability to manage AI-enhanced learning

### v2

@@ 5.3 A Step-by-Step Pedagogical Prompt Builder
To effectively scaffold students in creating pedagogical prompts, we divided the prompt construction process into six steps based on the six main components of a pedagogical prompt described in Section 3, according to the segmenting principle to reduce unnecessary mental effort (Mayer and Pilegard, 2005). For the Problem Context step, the builder lists key information that students must explicitly incorporate into their final prompt and presents it as a short-answer question. Other steps begin with a targeted multiple choice question (Figure 7-C1). After selecting or reading the corresponding step description, participants construct and submit a draft segment of the prompt (Figure 7-C2). When writing a prompt segment, students see their previously completed segments displayed before the current text input area as contextual reminders (Figure 6-Left). Students then submit their prompt segment to receive immediate, AI-generated feedback (Figure 7-C3). We defined specific acceptance criteria for each step of the prompt-building process, and employed OpenAI’s GPT-4o (specifically gpt-4o-2024-08-06) to provide personalized, elaborated actionable feedback, highlighting any omissions or areas for improvement. If the student cannot pass all the criteria after three attempts, a sample solution for this step will be provided (Figure 7-C4) to help the student finish the current prompting step (Figure 7-C5). Upon finishing constructing a prompt (Figure 6-Right), students are encouraged to use this prompt to interact with AI, such as putting it into their Chatbot of choice (e.g., ChatGPT, Claude) and observing the AI response.
@@
Figure 6: Left: Previous prompt segments. Students can view the segments they wrote in previous steps above the current input field. Right: Full prompt review. After completing the current scenario, students can view the full prompt built from prior steps.
6 User Study: Understand Student Learning and Perceptions
@@
Figure 7: Example of completing Problem Identification step in the prompt builder workflow. The student first (C1) considers and selects the most accurate option within the given list, and then reaches the (C2) prompt writing task, in which they are required to complete the current snippet of the prompt given their previous selection. When the student has drafted a prompt, (C3) an LLM-based feedback system will provide personalized feedback and provide the student with actionable suggestions. If the student still has difficulties passing the prompt validator, (C4) a sample solution will be shown to the student, and they have a chance to review and paste it into the answer box. (C5) When a student have successfully completed the prompt writing task for the current step, they will be directed to the next step.
@@
To understand the effectiveness of this learning intervention, we conducted a evaluation study with 22 students with IRB (Institutional Review Boards) approval. The RQs are:
@@
6.1 Participants & Study Procedure
@@
At the beginning of each session, participants first completed a consent with a pre-screening survey that included questions about their programming learning level, self-reported familiarity with AI and generative AI tools. Then they got to a pre-test that included three scenario-based prompt writing tasks (Figure 9) with a self-selected commercial AI chatbot, with ChatGPT as the default option. The scenarios were isomorphic across the pre-test, post-test, and learning session. After finishing all three tasks, participants provided reflections on prompts in their responses to help researchers gain a deeper understanding into their thought processes. The learning session started with a detailed on-boarding demo video. After that, participants completed three scenario-based learning tasks using our system in randomized order to decrease the ordering effect. Then they completed a post-test with isomorphic questions as the pre-test, followed by the reflection session. After the reflection, participants completed a post-survey containing 5-point Likert questions and open-ended feedback questions on their learning experiences and interaction with the system. All sections were not strictly time-limited, and participants were allowed to learn and finish the tests at their own pace. Each participant received a \$30 per hour gift card in local currency after the study. The detailed procedure can be found in Figure 8.
@@
Figure 8: User study procedure. Before the learning session: participants spent 25 minutes to complete a survey, 3 prompting questions in the pre-test, and reflect on the previous 3 prompts; learning session: participants began with watching a 2-3 minutes demo video of the system, and them spent 30-40 minutes to write pedagogical prompts for 3 different scenarios with scaffolding; after the learning session: participants completed 3 prompting questions in the post-test, reflected on the previous 3 prompts, and completed a post-survey in order.
@@
Figure 9: An example isomorphic scenario in pre-test with its corresponding prompt-writing task
6.2 Data Analysis
@@
To answer RQ1, two graders manually graded student written prompts in the pre-test and post-test using the rubric extracted from the detailed definition (Section 5.3.1) of a high-quality pedagogical prompt. Specifically, two graders independently graded 15% of the data, half randomly selected from the pre-test and half from the post-test. The graders reached high agreement in all fields (Cohen Kappa=0.72 - high agreement in inter-rater reliability (McHugh, 2012)) and further refined the grading rubric. With the refined rubric, one grader graded the remaining data. For RQ2, RQ3 and RQ4, we first summarized and reported students’ quantitative survey results, then uncovered the details based on their open-ended answers.
7 User Study Results
@@
7.1 RQ1: Students’ ability to write pedagogical prompts significantly improved after the learning intervention.
@@
For each component of a pedagogical prompt, we calculated students’ learning gain using post-test - pre-test. Then we applied Wilcoxon signed-rank tests (Gehan, 1965) because student test data is not normally distributed. Results show students’ (N=22) significant learning gains from pre-test to post-test for ALL 6 intended components (p $<$ .001 (see Table 1), indicating the initial success of this learning intervention in teaching pedagogical prompt.
@@
Figure 10: P5’s written prompt to the pre-test question and its corresponding isomorphic post-test question.
Table 1: Student test performance and learning gain for six pedagogical prompt components
@@
Tutoring Protocols 0.44 0.0 1.0 0.83 1.0 0.0 0.39 0.5 1.0 9.46 $<$.001*** 0.82
Difficulty Identification 0.70 1.0 1.0 0.95 1.0 0.0 0.26 0.0 0.8 9.91 $<$.001*** 0.86
AI’s Role 0.00 0.0 0.0 0.62 1.0 1.0 0.62 1.0 1.0 9.91 $<$.001*** 0.86
Problem Context 0.77 1.0 0.5 0.94 1.0 0.0 0.17 0.0 0.5 9.63 $<$.001*** 0.84
Learner’s Level 0.08 0.0 0.0 0.65 1.0 0.9 0.58 1.0 1.0 9.75 $<$.001*** 0.85
Guardrails 0.25 0.0 0.8 0.82 1.0 0.0 0.56 1.0 1.0 9.72 $<$.001*** 0.85
  • * $p$ $<$ .05, ** $p$ $<$ .01, *** $p$ $<$ .001. Wilcoxon signed-rank test used for non-parametric data; interquartile range (IQR) is reported to describe variability in scores.
7.2 RQ2: Students perceptions of their ability to effectively use LLM for learning have been corrected and increased after the learning intervention.
Students reported relatively high confidence in their ability to use LLMs effectively for learning in the pre-test (Figure 11, first row; M = 3.72 of 5). However, their self-reported confidence did not accurately reflect their actual prompting ability. Specifically, if students had an accurate self-assessment of their skills, a significant positive correlation between confidence and pre-test performance would be expected (Baxter and Norman, 2011). In contrast, in our result, a Pearson correlation showed a negative but not significant relationship between the perceived confidence score and pre-test performance (r = –.15, p = .498), suggesting a misalignment between the perception and actual abilities of the learners before the intervention. This pattern aligns with findings from Prather et al. (2024), suggesting that students may hold a false sense of confidence in their LLM-based learning, which can potentially lead to conceptual gaps and falling behind in course content. After the learning session, participants’ perceived confidence increased to an average of 4.63 of 5 (Figure 11-row 2), reflecting both greater perceived confidence from pre-survey to post-survey (p = .001, z = 5.38, effect size = .81, Wilcoxon signed-rank test) and a significant positive alignment between their perceived ability and actual performance (r = .54, p = .032, Pearson correlation). Furthermore, as indicated in Figure 11 (row 3 to row 4), students reported perceived improvements in identifying their specific struggles, although the improvement is not significant (p = .071, z = 5.26, Wilcoxon signed-rank test).
@@
Figure 11: Changes in students’ self-reported ability to use LLM tools for learning before and after the learning intervention
Figure 12: Students’ willingness to use pedagogical prompting in future learning and perceptions of the system experience reported in the post-survey
7.3 RQ3: Most students are extremely willing to apply pedagogical prompts in their future learning.
@@
7.4 RQ4: Students rated the system highly, despite occasional moments of frustration.
@@ 8 Discussion
8.1 Instructors: need for an interactive system that delivers brief instructional activities to teach pedagogical prompting for all CS novices
@@
8.1.1 Instructors prefer a brief but focused pedagogical prompting instruction session
@@
8.1.2 Pedagogical prompting should be taught to all students in CS classes, regardless of their major and level of programming.
There is a consensus on the targeted student population. Instructors think both CS majors and non-majors should learn pedagogical prompting, regardless of their major or level of programming. Recent studies support providing AI education to this broad audience (Shernoff, 2024; Walter, 2024; Denny et al., 2024). For instance, Walter (2024) emphasizes the importance of AI literacy and prompt engineering for all students regardless of their major programs, highlighting their role in fostering critical thinking skills essential in the age of AI (Walter, 2024). Similarly, Denny et al. (2024) discuss the transformative impact of generative AI on computing education, underscoring the need for students across disciplines to engage with these tools thoughtfully (Denny et al., 2024). These findings suggest that equipping all students with the skills to use AI tools pedagogically is crucial. It’s not only for students’ own academic success, but also for their preparedness in a workforce increasingly influenced by AI technologies.
@@
8.1.3 An interactive instructional system is the most preferred way for delivering the training
Regarding the medium of delivering pedagogical prompting instruction, an interactive system is the top choice. This finding demonstrates instructors’ preference for a more engaging instructional method. Their preferred ways align well with the ICAP framework (Chi and Wylie, 2014), which categorizes learning activities by the level of cognitive engagement they promote. Interactive instructional systems encourage students to either construct prompts themselves or actively select and refine them through system-guided scaffolding, both of which fall into the Constructive or Interactive modes of engagement. These levels are shown to produce deeper learning outcomes because learners generate new knowledge or meaning beyond the information presented. In contrast, alternatives such as demo videos or static spreadsheets primarily foster Passive or at best Active engagement, where students receive or manipulate information without generating new understanding. Practice questions can also be a viable option; however, compared to an interactive system, they may lack personalized, immediate feedback (Shute, 2008), as well as autonomy, which could induce boredom according to Pekrun and Perry (2014)’s Control-Value Theory of Achievement Emotions (Pekrun and Perry, 2014).
8.1.4 Selecting vs. Writing: Designing Interactions to Balance Efficiency and Active Skill Development
@@ 8.1.3 An interactive instructional system is the most preferred way for delivering the training
8.2 Learners: Developing Cognitive and Metacognitive Skills to Strengthen Learner Agency in the Age of Generative AI
@@
8.2.1 Prompting skill is a pressing need for nowadays learners, and scaffolded writing practice is an effective, immediate next-step that can significantly increase learners’ prompting ability in using AI for learning
@@
As generative AI becomes an inevitable component of future learning environments, it is crucial to prepare students to engage with it effectively rather than passively or improperly (Prather et al., 2025). Yet, learners, especially novices, often lack the strategies for effective AI-based help-seeking (Xiao et al., 2024a), and the misuses of AI can negatively impact learning outcomes (Bastani et al., 2024). Therefore, equipping students with the skills to use AI tools purposefully and critically will be an essential immediate next-step in education.
@@
Our system provides a solution to achieve this instructional goal in prompting learning in a short time period, as indicated by the promising results of the user study. After learning with our system, both learners’ ability and perceived confidence in pedagogical prompting significantly improved from pre-test to post-test. As shown in Table 1, participants made the most substantial gains in articulating the Learning Method and specifying the AI’s Role, which are critical components of effective pedagogical prompting. These improvements suggest that learners became more capable of providing structured, instructional guidance to the AI, thus enhancing the overall quality and clarity of their prompts. Mapping these gains to Nelson-Le Gall (1981)’s five-stage help-seeking model, we can observe strong developments in Stage 3: Identifying a potential helper, and Stage 4: Employing strategies to elicit help. By learning how to clearly specify the learning method and contextualize the AI’s role, students demonstrated an improved ability to recognize the AI as a valid and appropriate source of help for learning tasks. More importantly, they learned to strategically phrase their prompts in ways that actively engage the AI in a pedagogically useful manner.
@@
This success may be attributed to the system’s ability to stimulate learners’ psychological engagement (Stull and Mayer, 2007) through repeated practice (Pedersen et al., 2016) on constructive-level tasks (Chi and Wylie, 2014). We deliberately designed an interaction-heavy system, recognizing that while many existing works assume a high degree of learner self-regulation, this assumption often does not hold in real-world educational settings, as learners frequently demonstrate unregulated or suboptimal learning behaviors (Jansen et al., 2020; Molenaar et al., 2023; Xiao et al., 2024a). To address this, we embedded interactions that involve productive struggles (Gresalfi et al., 2009) and are difficult to "game," encouraging meaningful engagement without prerequisites on learners’ self-regulation. Since the effectiveness of the learning experience has been demonstrated, a next-step for future research is to investigate the system’s learning efficiency by optimizing learning gains while minimizing potential redundancy and learner frustration.
@@
8.2.2 Meta-cognitive skills can further strengthen students’ ability to manage AI-enhanced learning